# ReverseNER: A Self-Generated Example-Driven Framework for Zero-Shot Named Entity Recognition with Large Language Models

### v3

@@ 3.2.2 Example Label Generation
As shown in the dialogue above, the example labels generated through the reversed NER process may not always be semantically realistic, but they ensure reliability as valid example labels for NER tasks.
@@ 3.3 Example-Driven Inference
3.3.1 Example Labels Selection
To identify the most relevant example labels for each task sentence, a similarity matrix must be calculated between the sentences in the example library and those in the task set. To achieve this, we first use the BERT model to map each sentence in the example library into high-dimensional embedding vectors. Then, for each sentence in the task set, the cosine similarity between it and every sentence in the example library is calculated, resulting in a similarity matrix of size $N\times M$, where $N$ denotes the number of sentences in the task set, and $M$ denotes the number of sentences in the example library.
Based on the similarity matrix, we select the most relevant example labels to include in the prompt provided to the LLM during entity recognition.
@@ 3.3.2 Inference with LLMs
The constructed prompt to perform the NER task on a certain task sentence includes the following four elements: the entity type set to be recognized, definitions of each entity type, selected example labels, and the task sentence itself. The dialogue demonstration of inference is shown as follows:
Prompt Perform NER task on this entity type set: [PER, LOC, ORG]. (Omitted: Definitions of each entity type) Example 1: Input Sentence: John Smith and Marie Curie work for NASA in Paris. Output: [{"John Smith":"PER"}, {"Marie Curie":"PER"}, {"NASA":"ORG"}, {"Paris":"LOC"}] (Omitted: Example 2 to Example 5) Task Input Sentence: Lily and Mark collaborate at the World Health Organization in Geneva. Your output:
@@ 3.4 Entity-Level Self-Consistency Scoring
To mitigate the impact of LLM hallucinations, we perform multiple reasoning attempt rounds on the same sentence to be recognized, and introduce the self-consistency (SC) scoring mechanism. Different from the traditional Response-Level SC methods [20], we propose the Entity-Level SC Scoring method.
The occurrences of each inferred entity type across all reasoning attempts are counted, which is used as the self-consistency (SC) score for each entity. For each reasoning attempt, the score is calculated as the average SC score of all entities inferred in that response. The average SC score for each reasoning is formulated as:
$SC_{average}=\frac{1}{n}\sum_{i=1}^{n}SC(e_{i})$
where $n$ denotes the number of entities in the reasoning attempt, and $SC(e_{i})$ denotes the self-consistency score of entity $e_{i}$, which is the total number of times entity $e_{i}$ appears across all rounds of attempts.
Additionally, to handle cases where the correct output does not contain any entities, a majority-based rule is introduced: if more than half of the predictions indicate no entity, the final output is set to "no entity."
For each sentence in the task set, the self-consistency scoring method is applied during testing, and the result with the highest self-consistency score is selected as the final output.
@@ 4 Experiments
In this section, we experiment on four public benchmarks to evaluate the ReverseNER method in comparison to multiple baselines.
@@ 4.1 Setup
LLM
We utilize the GPT-4o mini model as the LLM to generate examples and perform the NER task to evaluate our algorithm. During the self-consistency scoring process, the temperature parameter is set to 0.8, and the prediction process is repeated five times.
Example Library Generation
For English texts, we utilize the BERT-base-uncased model [6] to derive sentence embeddings, while for Chinese texts, we use the BERT-base-Chinese model [6]. Sentences from the input set are clustered into 10 categories, and for each cluster’s central sentence i.e., the feature sentences, the LLM generates 3 examples by mimicking the sentence structure. Thus, these 30 example labels constitute the entire example library for a certain dataset.
Datasets
We use the following four datasets to evaluate the performance of the NER task. For English texts, we utilize the CoNLL03 dataset [18] and WikiGold dataset [2]. For Chinese texts, we use the People’s Daily dataset [21] and have constructed a novel dataset called GovAff, which consists of sentences pertaining to routine administrative affairs. The GovAff dataset encompasses four types of entities: person, government agency, corporation, and miscellaneous.
Baselines and Comparison
The performance achieved with the vanilla model, where no examples are generated or provided, and only the task sentence and entity types are provided to the model, is recorded as the baseline. Additionally, we conduct two few-shot NER experiments using 10 examples with gold labels and compare their results with the proposed method in this study.
@@ 4.2 Overall Performance
The results presented in Table 1 underscore the F1 score of the zero-shot ReverseNER method when compared to other methods. Notably, the zero-shot ReverseNER with self-consistency scoring attains the highest average F1 score of 79.10 across the four evaluated datasets, outperforming both the Vanilla zero-shot and few-shot baselines.
Relative to the Vanilla zero-shot baseline, the zero-shot ReverseNER with SC demonstrates a substantial enhancement in the average F1 score (79.10 vs. 70.97), reflecting a significant improvement in the model’s capacity to generalize to unseen data without the need for additional gold labels provided in advance. Furthermore, in comparison to the few-shot methods, both with and without SC, the zero-shot ReverseNER with SC exhibits superior performance, surpassing the few-shot SC approach (77.25 average) by 1.85 points.
Table 1: F1 Score of ReverseNER method and its comparisons.
Method CoNLL03 WikiGold People’s Daily GovAff Average
Vanilla 71.69 70.97 69.56 71.23 70.86
Few-shot, w/o SC 72.71 75.86 71.23 83.40 75.80
Few-shot, SC 74.21 76.76 73.89 84.14 77.25
Zero-shot ReverseNER, w/o SC 76.19 76.98 75.46 82.46 77.77
Zero-shot ReverseNER, SC 77.78 78.47 76.62 83.53 79.10
@@ 4.3.1 Detailed Performance
Table 2: Detailed NER Recognition Results. Methods with the best Precision for each metric are indicated with underlining, those with the best Recall are italicized, and those with the highest F1 score are highlighted in bold.
@@
Vanilla Precision 87.14 70.91 83.33 32.73 63.97
@@
F1 87.39 74.81 86.57 57.66 76.98
Zero-shot ReverseNER, SC Precision 92.16 76.88 86.89 54.39 77.66
Recall 85.65 73.30 85.12 70.15 79.29
F1 88.79 75.04 85.99 61.27 78.47
By examining the Precision, Recall, and F1 scores of each label across different methods, we observe that the ReverseNER with SC voting consistently achieves the highest performance in terms of Precision and F1 scores for most labels, followed by the ReverseNER without SC voting. The Vanilla method, however, frequently achieves higher Recall, particularly for PER and LOC types. This phenomenon can be attributed to two primary factors: (1) The conservative strategy employed by the LLM when generating examples, where uncertain terms that may not belong to the current entity label are often excluded, leads to higher Precision but potentially lower Recall; (2) The SC voting algorithm tends to penalize outputs containing entities that are not consistently present across all outputs, even if they are correct in some cases, which further boosts Precision while reducing Recall. Consequently, these strategies improve the F1 score, particularly for labels with less clearly defined boundaries such as MISC categories, as they achieve a better balance between Precision and Recall.
@@ 4.3.2 Impact of hyperparameters
Our ReverseNER method involves two hyperparameters: the number of clusters used during feature sentence extraction and the number of example labels that the LLM needs to construct per cluster. The product of these two hyperparameters determines the total number of labels in the example library. To evaluate the algorithm’s stability under different parameter settings, we do not enable the self-consistency scoring mechanism and run each hyperparameter combination 20 times, calculating the mean and standard deviation.
@@
Figure 2: F1 Score on WikiGold when the number of clusters is fixed at 10.
Table 3: F1 Score on WikiGold with different sets of hyperparameters, without self-consistency scoring
Clusters Labels per Cluster F1 Score
@@ 4.3.3 Effect of Entity-Level Self-Consistency Scoring
This isolated underperformance for the Entity-Level SC can be attributed to the fact that there are only three clearly defined types of entities in the People’s Daily Dataset: PER, LOC, and ORG, without the MISC type in other datasets, which does not have a well-defined boundary. The Entity-Level SC strategy emphasizes trading Recall for Precision to enhance the F1 score, but this approach is less effective in datasets where entity definitions are more straightforward, resulting in lower performance compared to the traditional Response-Level SC method.
Table 4: F1 Score comparison on different SC methods.
@@
WikiGold 78.47 78.00
People’s Daily 76.62 77.72

### v4

@@ 3.2.2 Entity-to-Sentence Expansion
As shown in the dialogue above, the entity-labeled examples generated through the reversed NER process may not always be semantically realistic. Still, the entities in the examples are correctly labeled, ensuring reliability as valid examples for NER tasks.
@@ 3.3 Example-Driven Inference
In this phase, the input consists of the full task set and the example library constructed in Section 3.2, while the output is the final NER result, with each task sentence annotated with its corresponding entities.
3.3.1 Example Selection
To identify the most relevant entity-labeled examples for each task sentence, a similarity matrix must be calculated between the sentences in the example library and those in the task set. To achieve this, we first use the BERT model to map each sentence in the example library into high-dimensional embedding vectors. Then, for each sentence in the task set, the cosine similarity between it and every sentence in the example library is calculated, resulting in a similarity matrix of size $N_{\mathcal{T}}\times N_{\mathcal{E}}$, where $N_{\mathcal{T}}$ denotes the number of sentences in the task set, and $N_{\mathcal{E}}$ denotes the number of sentences in the example library.
After constructing the similarity matrix, for each task sentence, we select the entity-labeled examples with the highest cosine similarity. Specifically, for a task sentence $t_{i}$ and an example sentence $e_{j}$, we find $K$ example sentences with the highest cosine similarity:
$\hat{\mathbf{e}}_{i}=\underset{\mathbf{e}_{j}\in\mathcal{E}}{\text{arg top-$K$}}\,\text{sim}(t_{i},e_{j}),$ (5)
where $t_{i}$ is the embedding vector of the $i$-th task sentence, $\mathbf{e}_{j}$ is the embedding vector of the $j$-th example sentence from the example library embedding set $\mathcal{E}$, and $\hat{\mathbf{e}}_{i}$ is the set of selected examples set for $t_{i}$. The arg top-$K$ operation selects $K$ example sentences with the highest cosine similarity for each task sentence.
Based on this, we select the most relevant entity-labeled examples to include in the prompt provided to the LLM during entity recognition.
@@ 3.3.2 Inference with LLMs
The constructed prompt to perform the NER task on a certain task sentence includes the following four elements: the entity type set to be recognized, definitions of each entity type, selected entity-labeled examples, and the task sentence itself. The LLM outputs the recognized entities within the task sentence. The dialogue demonstration of an inference attempt is shown as follows:
Prompt Perform named entity recognition task on this entity type set: [PER, LOC, ORG]. The definition of PER is …; The definition of LOC is …; The definition of ORG is …. Example 1: Input Sentence: John Smith and Marie Curie work for NASA in Paris. Output: [{"John Smith":"PER"}, {"Marie Curie":"PER"}, {"NASA":"ORG"}, {"Paris":"LOC"}] Example 2 to Example 5: … Task Input Sentence: Lily and Mark collaborate at the World Health Organization in Geneva. Your output:
@@ 3.4 Entity-Level Self-Consistency Scoring
To mitigate the impact of LLM hallucinations, we perform multiple inference attempts on each sentence and introduce a self-consistency (SC) scoring mechanism. Unlike traditional Response-Level SC methods [30], our Entity-Level SC Scoring method focuses on counting the occurrence of individual entities across all attempts. This process is optional, improving accuracy at the cost of increased computational resources.
We first count the occurrences of each recognized entity across all inference attempts, denoting this total count as the self-consistency score $SC(\mathbf{E}_{i})$ for entity $\mathbf{E}_{i}$. Next, for each inference attempt, we compute the average SC score for all entities predicted in that attempt. Formally, if an attempt predicts $n$ entities $\{\mathbf{E}_{1},\mathbf{E}_{2},\ldots,\mathbf{E}_{n}\}$, its self-consistency score $SC_{\text{attempt}}$ is formulated as:
$SC_{\text{attempt}}=\begin{cases}\displaystyle\frac{1}{n}\sum_{i=1}^{n}SC(\mathbf{E}_{i}),&\text{if }n>0,\\[10.00002pt] 0,&\text{if }n=0,\end{cases}$ (6)
In cases where most attempts predict at least an entity, we then select the single attempt with the highest $SC_{\text{attempt}}$ and use it as the final output. Additionally, to handle sentences that contain no entities, we apply a majority-based rule: if more than half of the attempts predict no entity, we label the sentence as “no entity.”
Entity-level SC is particularly beneficial when dealing with uncertain or ambiguous entities. Because each entity’s SC score depends on how frequently it appears across attempts, responses containing less certain entities (i.e., those that do not appear in all or most attempts) naturally receive lower average scores. Consequently, this approach sacrifices some recall (proportion of true entities identified) in favor of higher precision (proportion of predicted entities that are correct). Such precision-oriented selection is especially useful for categories like MISC in CoNLL03 and Wikigold, where the definitions of entities can be less clear or overlap with other labels.
@@ 4 Experiments
In this section, we experiment on four benchmarks to evaluate our ReverseNER method in comparison to multiple baselines.
@@ 4.1 Setup
LLM:
We utilize the GPT-4o mini model as the LLM backbone to generate examples and perform the NER task to evaluate our method. During the self-consistency scoring process, the temperature parameter is set to 0.8, and the inference attempt is repeated five times.
Example Library Construction:
For English texts, we utilize the BERT-base-uncased model [9] to derive sentence embeddings, while for Chinese texts, we use the BERT-base-Chinese model [9]. Sentences from the task set are clustered into 10 categories, and for each cluster’s central sentence i.e., the feature sentences, the LLM generates 3 examples by mimicking the sentence structure. Thus, these 30 entity-labeled examples constitute the entire example library for a certain dataset.
Datasets:
We use four datasets to evaluate the performance of the NER task. For English texts, we utilize the CoNLL03 dataset [26] and WikiGold dataset [4]. For Chinese texts, we use the People’s Daily dataset [31] and have constructed a novel dataset called GovAff, which consists of sentences pertaining to routine administrative affairs. The GovAff dataset encompasses four types of entities: person, government agency, corporation, and miscellaneous.
Baselines and Comparison:
The performance of the vanilla zero-shot method, which operates without any provided examples (relying solely on the task description, entity types, and their corresponding definitions—serves as the baseline for comparison). In addition, we perform a vanilla few-shot NER experiment using 10 pre-labeled examples and compare their results with those of the proposed method in this study.
Furthermore, we evaluate our approach against other zero-shot and few-shot methods leveraging LLMs, including the Self-improving Framework by Xie et al. [25] for zero-shot recognition and PromptNER [3] for both zero-shot and few-shot recognition. Given that these methods are tested on GPT-3.5, we also adopt GPT-3.5 as our LLM backbone for comparison. The results of the Self-improving Framework and PromptNER are directly taken from the best performances reported in their respective papers. Resource consumption is calculated based on the total number of LLM invocations, with $N_{\mathcal{T}}$ denoting the total number of sentences to be recognized in each task set.
@@ 4.2 Overall Performance
4.2.1 Comparison with Baseline Methods
The results presented in Table 1 demonstrate the superiority of the ReverseNER approach. On all datasets, ReverseNER without self-consistency, i.e., ReverseNER w/o SC, consistently outperforms the baseline methods, highlighting the method’s effectiveness in leveraging LLMs under zero-shot settings.
When incorporating self-consistency, ReverseNER with SC further boosts performance across all datasets. Notably, on WikiGold and GovAff, micro F1 of the method reaches 78.45 and 83.53, respectively, representing a significant improvement over baseline methods. A slight gain introduced by SC is also observed on People’s Daily, a dataset without ambiguous entities, where ReverseNER SC achieves 75.62 in micro F1, compared to 75.46 without SC.
Table 1: The micro F1 of ReverseNER method and its baseline comparisons with GPT-4o mini.
Method CoNLL03 WikiGold People’s Daily GovAff
Vanilla Zero-shot 71.69 70.97 67.93 78.10
Vanilla Few-shot 72.71 75.86 71.23 83.40
ReverseNER Zero-shot, w/o SC 76.19 76.98 75.46 82.46
ReverseNER Zero-shot, SC 77.78 78.45 75.62 83.53
4.2.2 Comparison with Existing Methods
Table 2 presents a comparison of the micro F1 achieved by our proposed ReverseNER method against existing zero-shot and few-shot approaches on the CoNLL03 and WikiGold datasets, along with an analysis of resource consumption. Compared to the Self-improving Framework with 1 Iteration, ReverseNER w/o SC demonstrates superior performance while significantly reducing resource consumption. Specifically, for low-resource scenarios, ReverseNER w/o SC attains micro F1 of 75.63 on CoNLL03 and 74.98 on WikiGold, utilizing only $10+10^{*}+N_{\mathcal{T}}$ LLM invocations. In contrast, the Self-improving Framework with 1 Iteration achieves 74.99 and 73.98 on CoNLL03 and WikiGold, respectively, but requires $500\times 5+N_{\mathcal{T}}$ LLM invocations.
Furthermore, for high-resource zero-shot methods, our ReverseNER with SC surpasses the Self-improving Framework with 8 Iterations on the WikiGold dataset, although it does not outperform it on the CoNLL03 dataset. Nevertheless, ReverseNER consistently exhibits lower resource consumption compared to the Self-improving Framework. Notably, the performance of ReverseNER with SC on the WikiGold dataset even exceeds the theoretical upper bound of the Self-improving Framework.
Table 2: The micro F1 and resource consumption of our ReverseNER method with its comparisons. All methods are evaluated using GPT-3.5. Resource consumption is measured by the total number of LLM invocations, where $N_{\mathcal{T}}$ represents the number of sentences in the task set. For the WikiGold dataset, $N_{\mathcal{T}}=1697$; for CoNLL03, $N_{\mathcal{T}}=3453$. Numbers or symbols marked with an asterisk (*) indicate that the output tokens required for this LLM invocation might significantly surpass the scope of a straightforward NER task.
Method CoNLL03 WikiGold Resource Consumption
Low-Resource Zero-shot Methods
Vanilla Zero-shot 68.98 70.77 $N_{\mathcal{T}}$
PromptNER, Zero-shot [3] 68.10 - $N_{\mathcal{T}}^{*}$
Self-improving Framework [25], 1 Iteration 74.99 73.98 $500\times 5+N_{\mathcal{T}}$
ReverseNER, w/o SC (Ours) 75.63 74.98 $10+10^{*}+N_{\mathcal{T}}$
High-Resource Zero-shot Methods
Self-improving Framework [25], 8 Iterations 77.49 73.98 $500\times 5\times 8+N_{\mathcal{T}}$
ReverseNER, SC (Ours) 76.77 76.72 $10+10^{*}+5N_{\mathcal{T}}$
Self-improving Framework [25], Theoretical Upper Bound 81.65 76.64 +$\infty$+ ∞
Few-shot Methods
Vanilla Few-shot 71.55 72.32 $N_{\mathcal{T}}$
PromptNER, Few-shot [3] 78.62 - $N_{\mathcal{T}}^{*}$
@@ 4.3 Analysis
4.3.1 Example Library Quality Evaluation
To enhance the interpretability of our ReverseNER method, we use the micro F1 metric to evaluate the correctness of entity labels across the entire example library. We also define two additional metrics to assess the quality of our example library: Average of Highest Cosine Similarities (AHS) and Entity Diversity Ratio (EDR), focusing on similarity and diversity, respectively. Below are the definitions and explanations for each metric:
Micro F1: The micro F1 assesses the overall performance across all entity labels within the example library. It ranges from 00 to $100$, with higher values indicating better performance. Since an LLM entirely generates the example library, this metric is obtained through manual inspection to identify any missing entities or labeling errors.
Average of Highest Cosine Similarities (AHS): AHS measures the overall alignment between the input task set and the example library by averaging the highest cosine similarity scores for each sentence in the input dataset. It ranges from 00 to $1$, where higher values signify greater similarity. Specifically, given a task set $\mathcal{T}$ containing $N_{\mathcal{T}}$ sentences $\{t_{1},t_{2},...t_{N_{\mathcal{T}}}\}$ and an example library $\mathcal{E}$, AHS is calculated as follows:
$\overline{\text{sim}}(t_{i})=\max_{e_{j}\in\mathcal{E}}\,\text{sim}(t_{i},e_{j}),\$
$\text{AHS}=\frac{1}{N_{\mathcal{T}}}\sum_{i=1}^{N_{\mathcal{T}}}\overline{\text{sim}}(t_{i}).$ (7)
This metric quantifies the overall alignment between the input and example datasets by assessing how closely each input sentence matches its most similar counterpart in the example library.
Entity Diversity Ratio (EDR): EDR compares the semantic diversity of entities in the example library against that of the input dataset. It is defined for each entity type $\tau$ and ranges from 00 to $\infty$, with values closer to $1$ indicating similar levels of diversity. The Entity Diversity (ED) for a given entity type of the example library or task set is defined as:
[…]
where $\mathbf{E}_{i}$ and $\mathbf{E}_{j}$ are two unique entities of the same type $\tau$ within the same example library or task set, and $N_{\tau}$ is the total number of unique entities for that type.
The Entity Diversity Ratio (EDR) for entity type $\tau$ is then computed as the division of ED of the example library $\text{ED}_{\mathcal{E}}$ and the ED of the task set $\text{ED}_{\mathcal{T}}$:
[…]
An EDR greater than 1 indicates that the example library has greater semantic diversity for the entity type compared to the task set, an EDR less than 1 means the task set has greater semantic diversity, and an EDR equal to 1 signifies identical levels of semantic diversity between the example library and the task set for that entity type.
Note that EDR is only about diversity. An EDR value being close to $1$ does not necessarily imply that the entities in the example library cover the same scope as those in the task set. Therefore, the EDR metric should be considered in conjunction with the AHS index for a more comprehensive evaluation.
The metrics of the constructed example library for each dataset in Section 4.1 with GPT-4o mini are presented in Table 3. A key achievement is the correctness of the constructed example libraries, which all achieved a perfect micro F1 of 100.00. Given that each example library contains only 30 entity-labeled sentences, while the dataset contains thousands of sentences, the AHS metric values, ranging from 0.6934 to 0.7737, remain promising. Although the EDR metric is close to 1 for many entities, it shows less promise for some entities, such as MISC in the CoNLL03-based example library and GOV in the GovAff-based example library. The definitions of these entities are often ambiguous and encompass a broad semantic range.
Table 3: Quality evaluation of constructed example library by correctness (micro F1), similarity to the dataset (AHS), and entity diversity ratio (EDR). Base Dataset means the example library is built for the certain dataset.
Base Dataset Micro F1 AHS EDR 1 EDR 2 EDR 3 EDR 4
CoNLL03 100.00 0.7737 PER: 1.5956 ORG: 1.0517 LOC: 0.8346 MISC: 0.7011
WikiGold 100.00 0.7664 PER: 1.5425 ORG: 1.0784 LOC: 0.9556 MISC: 0.9822
People’s Daily 100.00 0.6934 PER: 1.3497 ORG: 1.0699 LOC: 1.3835 -
GovAff 100.00 0.7107 PER: 1.2124 CORP: 0.9389 GOV: 0.6900 MISC: 0.7674
@@ 4.3.2 Detailed Performance
Table 4: Detailed NER results on WikiGold dataset with GPT-4o mini. Methods with the best Precision for each metric are indicated with underlining, those with the best Recall are italicized, and those with the highest F1 are highlighted in bold.
@@
Vanilla Zero-shot Precision 87.14 70.91 83.33 32.73 63.97
@@
F1 87.55 74.81 86.57 57.66 76.98
Zero-shot ReverseNER, SC Precision 91.26 80.78 85.69 55.53 78.43
Recall 84.11 70.68 86.03 70.15 78.45
F1 87.54 75.39 85.86 61.99 78.45
By examining the Precision, Recall, and F1 scores of each label across different methods, we observe that the ReverseNER with SC voting consistently achieves the highest performance in terms of Precision and F1 scores for most labels, followed by the ReverseNER without SC voting. The vanilla zero-shot method, however, frequently achieves higher Recall, particularly for PER and ORG types. This phenomenon can be attributed to two primary factors: (1) The conservative strategy employed by the LLM when generating examples, where uncertain terms that may not belong to the current entity label are often excluded, leads to higher Precision but potentially lower Recall; (2) The SC voting algorithm tends to penalize outputs containing entities that are not consistently present across all outputs, even if they are correct in some cases, which further boosts Precision while reducing Recall. Consequently, these strategies improve the F1 score, particularly for labels with less clearly defined boundaries such as MISC categories, as they achieve a better balance between Precision and Recall.
@@ 4.3.3 Impact of hyperparameters
Our ReverseNER method involves two hyperparameters: the number of clusters used during feature sentence extraction and the number of entity-labeled examples that the LLM needs to construct per cluster. The product of these two hyperparameters determines the total number of labels in the example library. To evaluate the algorithm’s stability under different parameter settings, we do not enable the self-consistency scoring mechanism and run each hyperparameter combination 20 times with GPT-4o mini, calculating the mean and standard deviation.
@@
Figure 3: Average micro F1 on WikiGold when the number of clusters is fixed at 10, evaluated on GPT-4o mini.
Table 5: The micro F1 on WikiGold with different sets of hyperparameters, without self-consistency scoring, evaluated on GPT-4o mini.
Clusters Labels per Cluster Micro F1
@@ 4.3.4 Effect of Entity-Level Self-Consistency Scoring
This isolated underperformance for the Entity-Level SC can be attributed to the fact that there are only three clearly defined types of entities in the People’s Daily Dataset: PER, LOC, and ORG, without the MISC type in other datasets, which does not have a well-defined boundary. The Entity-Level SC strategy textslasizes trading Recall for Precision to enhance the F1 score, but this approach is less effective in datasets where entity definitions are more straightforward, resulting in lower performance compared to the traditional Response-Level SC method.
Table 6: The micro F1 comparison on different SC methods, evaluated on GPT-4o mini.
@@
WikiGold 78.45 78.00
People’s Daily 75.62 77.72
@@
4.3.5 Performance on Lightweight LLMs
The performance of our method, evaluated on Qwen2.5-7b (qwen2.5-7b-instruct) [19], is shown in Table 7. Our method provides a noticeable improvement in the zero-shot NER performance of Qwen2.5-7b, especially for the Chinese datasets, People’s Daily and GovAff. This improvement can be attributed to the better performance of the Qwen series models on Chinese text and the higher quality of the constructed example library. Notably, the performance of Qwen2.5-7b on the People’s Daily dataset is somewhat disappointing. One reason for this is that many sentences in the dataset triggered its safety mechanism, causing the LLM to refuse to provide any responses.
Table 7: The micro F1 comparison on Qwen2.5-7b.
Dataset Vanilla Zero-shot ReverseNER, w/o SC
CoNLL03 60.36 62.46
WikiGold 63.88 66.70
People’s Daily 52.37 57.02
GovAff 72.50 77.05
4.3.6 Handling Isolated Zero-shot NER
We define Isolated Zero-shot NER as the scenario in which each sentence is treated as an independent unit. In this setting, the model processes one sentence at a time, ensuring that only the information within that sentence is available for entity recognition. This contrasts with previous approaches where multiple sentences or entire documents might be processed together, potentially allowing context from surrounding sentences to influence the NER results.
Our ReverseNER can handle this challenge without any modifications. By treating each task sentence as a complete task set, the feature sentence extraction process in Section 3.1 reduces to selecting the sentence itself as the feature sentence. We then construct the example library by mimicking its structure and obtain 3 similar entity-labeled sentences as examples. Finally, we use these examples to guide the LLM during inference.
Taking WikiGold as a case study, we emulate an Isolated Zero-shot NER scenario on it. The results and computational resource consumption are shown in Table 8. In the Isolated Zero-shot NER setting, our ReverseNER algorithm still demonstrates significant improvements over the baseline. In contrast, all other zero-shot methods used for comparison in previous sections simply cannot perform. However, when performing Isolated Zero-shot NER with our method, the need to construct an example library for each sentence increases the computational resource consumption.
Table 8: The micro F1 and resource consumption of Isolated Zero-shot NER on WikiGold with GPT-4o mini. Resource consumption is measured by the total number of LLM invocations, $N_{\mathcal{T}}=1697$. Symbols marked with an asterisk (*) indicate that the output tokens required for this LLM invocation might significantly surpass the scope of a straightforward NER task.
Method micro F1 Resource Consumption
Vanilla Zero-shot 70.97 $N_{\mathcal{T}}$
ReverseNER, w/o SC 76.22 $2N_{\mathcal{T}}+N_{\mathcal{T}}^{*}$
ReverseNER, SC 77.77 $6N_{\mathcal{T}}+N_{\mathcal{T}}^{*}$